Evaluate:

  1. &Integral 3 (x2 -4x + 4x – 2) dx
1
Expert's answer
2021-03-16T08:44:11-0400

"\int(x² - 4x + 4x^{-2}) dx =\\
\int(x² -4x + \dfrac{4}{x²}) dx = \\
\int x²\ dx - 4\int x \ dx+4 \int \dfrac{1}{x²}\ dx =\\
\dfrac{x³}3 - 4(\dfrac{x²}2) + 4(-\dfrac1x)+c ="


"\dfrac{1}3 x^3 - 2x^2 - 4x^{-1}+c"